Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-2021-05-11

Observers: Daniele Bjørn Malesani, Giorgos Leloudas
Support: Miika Pursiainen, Panos Charalampopoulos
TAT: Giuliano Pignata, Santiago Gonzalez, Jacob Teffs

EFOSC Calibrations: 
Bias, dome_flats, arcs taken in the afternoon.
Bias, dome_flats, arcs taken in the afternoon.


Weather:
Start: Humidity ~20 %, wind ~ 5 m/s
- clear sky.
Middle:
Humidity 20 % , wind ~5 m/s , clear
End:

Comments:
- The telescope operator reports that they found a workaround to the problem that affected the guide probe last night.
- Seeing in the beginning of the night quite poor >2"; for this reason we skipped some faint targets in favor of brighter ones.


 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
EFOSC2
        
22:36
Telescope opens






Well in twilight, preset to std and check
22:36L745aStandardg13, g11, g161"
Various
1.1
 Everything seems OK
22:46
SN2021dov
Follow-up
g11, g16
1"
1500 s + 1500 s
1.1
1.4"
Started very early in twilight (23:23 UT)
Quick-look of the spectra show good S/N for g11
But maybe the object moved out of the slit for g16 ?!
23:54
ZTF21aauufbz
Classification
g13
1"
900 s
1.8
2.0"
Seeing getting bad, S/N not excellent.
Possible TDE but mostly galaxy.
00:22
SDSSJ0952+2143
Follow-up
g13
1.5"
2700 s
1.8
2.2"
Because of bad seeing, changed from g11+16 to g13
01:17
SN2021hpx
Follow-up
g11
1.5"
1800 s
1.2
1.9"

01:54
SN2021gno
Follow-up
g13
1.5"
2700 s
1.4
1.7"
Seeing slightly improving
02:47
AT2021lei
Classification
g13
1.0"
1500 s
1.4
1.6"
Some sort of Ic-BL. Host lines at z = 0.112 ?
03:18
SN2020acat
Follow-up
g11/g16
1.0"
 2 x 1800 s
1.2
1.5"

04:42
SN2021kgj
Follow-upg11/g161.0"
 2 x 1500 s
1.5
1.4"

05:55
LTT7379
Standardg13, g11, g161", 1.5"
Various
1.1
1.3"

06:20    
SN2021fpl
Follow-upg11/g161.0"
 2 x 1800 s
1.4
1.4"
Problem with calibration unit. Skipping flats for g16.
07:35
SN2021een
Follow-upg13
1.0", 1.5"
 2 x 2700 + 1 x 1500 s
1.7
1.7" - 2.1"
 The seeing was 1.7" at the start of the 1st exposure became 2.1" at the second and 1.8" at the third. We therefore changed the slit to 1.5" for the second and 3rd exposure. The last exposure was stopped at 1500 sec because of twilight but data is saved.
10.10
EG131
Standardg11, g13, g16
1", 1.5" Various
1.2"




















Comments